OpenStudy (ms2016):

A 33 gram sample of a substance that's used to detect explosives has a k-value of 0.1473. What is the substance's half-life in days, rounded to the nearest tenth?

OpenStudy (wolf1728):

half-life = ln(2) / k half-life = .693/0.1473 half-life = 4.7 would that be days ?

OpenStudy (wolf1728):

Does the k-value have any units?

OpenStudy (ms2016):

Yes, Thanks for the help!